OpenStudy (jane11509):

Which pair of numbers have a GCF of 7? Choose all answers that are correct. A. 21 and 3 B. 7 and 14 C. 7 and 1 D. 28 and 7

OpenStudy (mathmath333):

find the common factors of all terms

OpenStudy (mathmath333):

a.) 21=7*3 3=3*1 ,here gcf =3

OpenStudy (jane11509):

Can you break it down a little further please?

OpenStudy (anonymous):

C

OpenStudy (jane11509):

thank you ^_^